# Learning to Play Two-Player Perfect-Information Games without Knowledge

### v4

@@ 5.2 Completion
It is not clear, however, that this improves performance as it prunes a portion of the game tree in the exploration whose values could be useful for learning (but could also negatively impact the learning performance).
@@ 5.3 Comparison of Reinforcement Heuristics
We now compare the different heuristics that we have proposed to the classical terminal evaluation function $\mathrm{b_{t}}$ on different games, using the protocol of Section 3.2. Each combination uses descent with completion (Algorithm 9) and completed $\epsilon$-greedy (see Algorithm 2 and Section 5.2). Each combination uses a different terminal evaluation function. These terminal evaluations are the classical (“binary”) evaluation function $\mathrm{b_{t}}$, the additive depth heuristic, the multiplicative depth heuristic, the scoring heuristic, the cummulative mobility, and the presence heuristic. Other parameters are the same as Section 3.3. There are, at most, a total of $6$ combinations per game (on some games, some heuristics are not evaluated because they are trivially of no interest or equivalent to another heuristic). The experiment was repeated $48$ times. The winning percentage of a combination for each game and for each evaluation step (i.e. each hour) is therefore calculated from $96$ to $192$ matches. The final winning percentages are shown in Table 5. Each percentage of the table has required between $4608$ and $9216$ matches. On average and in $7$ of the $9$ games, the classic terminal heuristic has the worst percentage (exception are Othello and Lines of Action). In scoring games, scoring is the best heuristic, as we might expect. Leaving aside the score heuristic, with the exception of Surakarta, Othello and Clobber, it is one of the two depth heuristics that has the best winning percentage. In Surakarta and Clobber, mobility is just ahead of the depth heuristics. On average, using the additive depth heuristic instead of using the classic evaluation increases the winning percentage by $15\%$, and using the best depth heuristic increases the winning percentage by $19\%$. The winning percentage curves are shown in Figure 5. The final percentages summarize the curves quite well. Note however, on the one hand, that the clear impact compared to the other heuristics (except score) of the additive depth heuristic to Breakthrough, Amazons, Hex, and Santorini and of the multiplicative depth heuristic to Clobber, Hex, and Open Outer Gomoku. In conclusion, the use of generic reinforcement heuristics has significantly improved performances and the depth heuristics are prime candidates as a powerful generic reinforcement heuristic.
@@
Table 5: Final winning percentages of the combinations of the experiment of Section 5.3 (X: heuristic without interest in this context ; presence coincides with score in Surakarta and Othello ; see 3.2 ; $95\%$ confidence intervals: max ±$\pm 1.04\%$)
6 Search Algorithms for Game Playing
In this section, we propose another variant of $\mathrm{UBFM}$, dedicated to be used in competition mode. Then, we compare it with other tree search algorithms.
7 Ordinal Distribution and Application to Hex
In this section, we propose the last technique, a new action selection distribution, and we apply it with all the previous techniques to design program-players to the game of Hex.
7.1 Ordinal Action Distribution
Thus, we propose an alternative probability distribution (see Section 2.3), that we call ordinal distribution. This distribution does not depend on the value of states. However, it depends on the order of their values. Its formula is:
@@ 7.1 Ordinal Action Distribution
with $n$ the number of children of the root, $i\in\{0,\ldots,n-1\}$, $c_{i}$ the $i$-th best child of the root, $P\left(c_{i}\right)$ the probability of playing the action leading to the child $c_{i}$ and $\epsilon^{\prime}$ the exploitation parameter ($\epsilon^{\prime}=1-\epsilon$). Algorithm 11 describes the action selection method resulting from the use of the ordinal distribution with an optimized calculation.
Function ordinal($s$, $v^{\prime}$)
@@
            $A\leftarrow\text{{{actions(}}{\emph{\text{s}}}{{)}}}$ sorted in descending order by $a\mapsto v^{\prime}(s,a)$
@@
            $A\leftarrow\text{{{actions(}}{\emph{\text{s}}}{{)}}}$ sorted in ascending order by $a\mapsto v^{\prime}(s,a)$
      $j\leftarrow 0$
@@
            if probability $\left(\frac{t}{T}\cdot\left(n-j-1\right)+1\right)/\left(n-j\right)$  then
@@
            $j\leftarrow j+1$
Algorithm 11 Ordinal action distribution algorithm with simulated annealing used in the experiments of this article (see Table 1 for the definitions of symbols).

### v5

@@ 5.2 Completed Algorithms
It is not clear, however, that completed action selection improves, in practice, performance as it prunes a portion of the game tree whose values could be useful for learning (but it would be surprising if this were not the case).
Function ubfms_iteration($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$)
      if terminal($s$) then
            $S\leftarrow S\cup\{s\}$
            $r(s),c(s),v(s)\leftarrow 1,\mathrm{b_{t}}(s),f_{\mathrm{t}}(s)$
      else
            if $r\left(s\right)=0$ then
                  if $s\notin S$ then
                        $S\leftarrow S\cup\{s\}$
                        foreach $a\in$ actions($s$) do
                              if terminal($a\left(s\right)$) then
                                    $r(s,a),c(s,a),v(s,a)\leftarrow$ ubfms_iteration($a(s),S,T,f_{\theta},f_{\mathrm{t}}$)
                              else
                                    $v(s,a)\leftarrow f_{\theta}\left(a(s)\right)$
                  else
                        $A\leftarrow\left\{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}\,|\,r\left(s,a\right)=0\right\}$
                        $a\leftarrow$ completed_best_action_dual($s$, $A$)
                        $n(s,a)\leftarrow n(s,a)+1$
                        $r\left(s,a\right),c\left(s,a\right),v\left(s,a\right)\leftarrow$ ubfms_iteration($a(s),S,T,f_{\theta},f_{\mathrm{t}}$)
                  $a\leftarrow$ completed_best_action($s$, actions(s))
                  $c(s),v(s)\leftarrow c\left(s,a\right),v\left(s,a\right)$
                  $r\left(s\right)\leftarrow$ backup_resolution($s$)
      return $r(s),c(s),v(s)$
Function ubfms_tree_search($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$, $\tau$)
      $t=$ time()
      while time()$-\,t<\tau\wedge r\left(s\right)=0$ do ubfms_iteration($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$)
      return $S$, $T$
Algorithm 12 $\mathrm{UBFM}_{\mathrm{s}}$ tree search algorithm with completion and resolution stop (see Table 1 for the definitions of symbols and Algorithm 11 for the definitions of completed_best_action($s$) and backup_resolution($s$)). Note: $T=(v,c,r,n)$. Note: Adding terminal states in $S$ is only useful during training with tree learning, so it should not be done during confrontations.
Function safest_action($s$, $T$)
      if first_player($s$) then
            return ${\displaystyle\operatorname*{\mathrm{arg\,max}}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}\left(c(s,a),n(s,a),v\left(s,a\right)\right)}$
      else
            return ${\displaystyle\operatorname*{\mathrm{arg\,min}}_{a\in\textnormal{{actions(}}\textnormal{\emph{s}}\textnormal{{)}}}\left(c(s,a),-n(s,a),v\left(s,a\right)\right)}$
Function ubfms($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$, $\tau$)
      $S,T\leftarrow$ubfms_tree_search($s$, $S$, $T$, $f_{\theta}$, $f_{\mathrm{t}}$,$\tau$)
      return safest_action($s$, $T$)
Algorithm 13 $\mathrm{UBFM}_{\mathrm{s}}$ action decision algorithm with completion (see Algorithm 12 for the definition of the method ubfms_tree_search() ; see Table 1 for the definitions of symbols). Note: $T=(v,c,r,n)$ and $S$ is the set of states of the game tree which are non-leaves or terminal.
Remark 10.
In some application cases, we will prefer to ensure the draw rather than trying to win. Algorithm 13 must then be adapted to decide the action to play. If the first player prefers to guarantee draws, it must instead maximize:
$\left(c(s,a),r\left(s,a\right),n(s,a),v\left(s,a\right)\right)\text{.}$
If the second player prefers to guarantee draws, it must instead minimize:
$\left(c(s,a),-r\left(s,a\right),n(s,a),v\left(s,a\right)\right)\text{.}$
Remark 11.
Remark 12.
For games without draw, the completion of Unbounded Minimax can be simply implemented by giving a value of +$\infty$+ ∞ to winning terminal states and a value of $\infty$- ∞ to losing terminal states (that is to say, it is sufficient to replace the terminal function by this infinite terminal function).
However, in the additional special case of adaptive evaluation functions with values in ]1,1[,]-1,1[,] - 1 , 1 [ , it is not sufficient that the terminal evaluation functions are in $\left\{-1,1\right\}$. Indeed, in practice, with rounding errors, the adaptive evaluation function being real value will have values in $[-1,1]$.
6 Reinforcement Heuristic to Improve Learning Performance
In this section, we propose the technique of reinforcement heuristic, which consists to replace the classical terminal evaluation function – that we denote by $\mathrm{b_{t}}$, which returns $1$ if the first player wins, $-1$ if the second player wins, and 00 in case of a draw [38, 2, 14] – by another heuristic to evaluate terminal states during the learning process. By using the reinforcement heuristic technique, non-terminal states are therefore evaluated differently, partial game trees and thus matches during the learning process are different, which can impact the learning performances. We start by defining what we call a reinforcement heuristic and we offer several reinforcement heuristics. Finally, we compare the reinforcement heuristics that we propose to the classical terminal evaluation function.
6.1 Reinforcement Heuristic Definition
Definition 13.
6.2 Some Reinforcement Heuristics
6.2.1 Scoring
Some games have a natural reinforcement heuristic: the game score. For example, in the case of the game Othello (and in the case of the game Surakarta), the game score is the number of its pieces minus the number of pieces of his opponent (the goal of the game is to have more pieces than its opponent at the end of the game). The scoring heuristic used as a reinforcement heuristic consists of evaluating the terminal states by the final score of the game. With that reinforcement heuristic, the adaptive evaluation function will seek to learn the score of states. In the context of an algorithm based on minimax, the score of a non-terminal state is the minimax value of the subtree starting from that state whose terminal leaves are evaluated by their scores. After training, the adaptive evaluation function then contains more information than just an approximation of the result of the game, it contains an approximation of the score of the game. If the game score is intuitive, the scoring heuristic should improve learning performances.
Remark 14.
In the context of the game of the Amazons, the score is the size of the territory of the winning player, i.e. the squares which can be reached by a piece of the winning player. It is approximately the number of empty squares.
6.2.2 Additive and Multiplicative Depth Heuristics
Now we offer the following reinforcement heuristic: the depth heuristic. It consists in giving a better value to the winning states close to the start of the game than to the winning states far from the start. Reinforcement learning with the depth heuristic is learning the duration of matches in addition to their results. This learned information is then used to try to win as quickly as possible and try to lose as late as possible. If learning durations can be done, this should reduce the length of games and therefore speed up learning (by increasing the number of matches played in the same time but also by reducing the propagation time of end-of-game information to the start-of-game states). Seeking to lose for as long as possible helps avoid missing out on potential long-term wins (even if it makes games last longer).
We propose two realizations of the depth heuristic: the additive depth heuristic, that we denote by $\mathrm{p_{t}}$, and the multiplicative depth heuristic, that we denote by $\mathrm{p_{t}}^{\prime}$. The evaluation function $\mathrm{p_{t}}$ returns the value $l$ if the first player wins, the value $-l$ if the second player wins, and 00 in case of a draw, with $l=P-p+1$ where $P$ is the maximum number of playable actions in a game and $p$ is the number of actions played since the beginning of the game. For the game of Hex, $l$ is the number of empty cells on the board plus $1$. For the games where $P$ is very large or difficult to compute, we can instead use $l=\max\left(1,\tilde{P}-p\right)$ with $\tilde{P}$ a constant approximating $P$ (close to the empirical maximum length of matches). The evaluation function $\mathrm{p_{t}}^{\prime}$ is identical except that $l$ satisfies $l=\frac{P^{\prime}}{p}$, and $P^{\prime}$ can be the exact or empirical average length of matches.
Remark 15.
6.2.3 Cummulative Mobility
The next reinforcement heuristic that we propose is cummulative mobility. It consists in favoring the matchs where the player has more possibility of action and where his opponent has less. The implementation used in this article is as following. The value of a terminal state is $\frac{M_{1}}{M_{2}}$ if the first player wins, $-\frac{M_{2}}{M_{1}}$ if the second player wins, and 00 in case of a draw, where $M_{1}$ is the mean of the number of available actions in each turn of the first player since the start of the game and $M_{2}$ is the mean of the number of available actions in each turn of the second player since the start of the game.
6.2.4 Piece Counting: Presence
Finally, we propose as reinforcement heuristic: the presence heuristic. It consists in taking into account the number of pieces of each player. This heuristic is based on the principle that the more a player has pieces the more this one has an advantage. There are several implementations for the presence heuristic. We use in this article the following implementation: the heuristic value is $\max(n_{1}-n_{2},1)$ if the first player wins, $\min(n_{1}-n_{2},-1)$ if the second player wins, and 00 in case of a draw, where $n_{1}$ is the number of pieces of the first player and $n_{2}$ is the number of pieces of the second player. Note that in the games Surakarta and Othello, the score corresponds to a presence heuristic.
@@ 6.3 Comparison of Reinforcement Heuristics
We now compare the different heuristics that we have proposed to the classical terminal evaluation function $\mathrm{b_{t}}$ on different games, using the protocol of Section 3.3. Each combination uses Descent with completion (Algorithm 10), completed $\epsilon$-greedy (see Algorithm 2, and Section 5). Each combination uses a different terminal evaluation function. These terminal evaluations are the classical game gain evaluation function $\mathrm{b_{t}}$, the additive depth heuristic, the multiplicative depth heuristic, the scoring heuristic, the cummulative mobility, and the presence heuristic. Other parameters are the same as Section 3.4. There are, at most, a total of $6$ combinations per game (on some games, some heuristics are not evaluated because they are trivially of no interest or equivalent to another heuristic). The experiment was repeated $48$ times. The winning percentage of a combination for each game and for each evaluation step (i.e. each hour) is therefore calculated from $288$ to $576$ matches.
The final winning percentages are shown in Table 4. Each percentage of the table has required between $13,824$ and $27,648$ matches. On average and for $7$ of the $9$ games, the classic terminal heuristic has the worst percentage (exception are Othello and Lines of Action). In scoring games, scoring is the best heuristic, as we might expect. Leaving aside the score heuristic, with the exception of Surakarta, Othello and Clobber, it is one of the two depth heuristics that has the best winning percentage. In Surakarta and Clobber, mobility is just ahead of the depth heuristics. On average, using the additive depth heuristic instead of using the classic evaluation increases the winning percentage by $15\%$, and using the best depth heuristic increases the winning percentage by $19\%$. The winning percentage curves are shown in Figure 4. The final percentages summarizes the curves quite well. Note that the positive impact of the additive depth heuristic compared to the other heuristics (except score) is particulary clear at Breakthrough, Amazons, Hex, and Santorini. In a similar manner, the positive impact of the multiplicative depth heuristic is particulary clear at Clobber, Hex, and Open Outer Gomoku.
In conclusion, the use of generic reinforcement heuristics has significantly improved performances and the depth heuristics are prime candidates as a powerful generic reinforcement heuristic.
@@
Table 4: Final winning percentages of the combinations of the experiment of Section 6.3 (X: heuristic without interest in this experiment or for the associated game ; presence coincides with score in Surakarta and Othello ; $95\%$ confidence intervals: max ±$\pm 1.04\%$)
7 Ordinal Distribution
In this section, we propose another technique, a new action selection distribution. It is based on a new, alternative, probability distribution, that we call ordinal distribution.
The ordinal distribution does not depend on the value of states. However, it depends on the order of their values. On the one hand, this distribution has the advantage of providing different probability distributions for actions that are not the best, unlike $\epsilon$-greedy. On the other hand, it has a more intuitive interpretation than the softmax function. Indeed, with this distribution, we choose with probability $p$ to play the best action. If we do not choose it, we choose the second best action with probability approximately $p$. Otherwise, we choose the third action with probability approximately $p$, etc. More precisely, with this distribution, the conditional probability of playing an action knowing that we do not play a better action is a linear interpolation between the uniform distribution and the max distribution (with respect to the exploitation parameter $\epsilon^{\prime}$). Its formula is:
@@ 7 Ordinal Distribution
with $n$ the number of children of the root, $i\in\{0,\ldots,n-1\}$, $c_{i}$ the $i$-th best child of the root, $P\left(c_{i}\right)$ the probability of playing the action leading to the child $c_{i}$, and $\epsilon^{\prime}$ the exploitation parameter ($\epsilon^{\prime}=1-\epsilon$). Algorithm 14 describes the action selection method resulting from the use of the ordinal distribution with an optimized calculation.
Function ordinal($s$, $v$)
@@
            $A\leftarrow\text{{{actions(}}{\emph{\text{s}}}{{)}}}$ sorted in descending order by $a\mapsto v(s,a)$
@@
            $A\leftarrow\text{{{actions(}}{\emph{\text{s}}}{{)}}}$ sorted in ascending order by $a\mapsto v(s,a)$
      $i\leftarrow 0$
@@
            if  $\mathrm{random}\left(\right)\leq\left(\frac{t}{T}\cdot\left(n-i-1\right)+1\right)/\left(n-i\right)$  then
@@
            $i\leftarrow i+1$
Algorithm 14 Ordinal action distribution algorithm with simulated annealing ($\epsilon^{\prime}=\frac{t}{T}$) used in the experiments of this article (see Table 1 for the definitions of symbols).